### v1

# Bifurcations in monotonic viscous boundary layers

Abstract
It is well-established that shear layer flows are linearly unstable provided the viscosity is small enough, when the horizontal Fourier number lies in some interval, between the so called lower and upper marginally stable curves. In this article, we prove that monotonic shear layers undergo a bifurcation near the upper marginally stable curve. In the particular case of the exponential boundary layer, we give evidences that this bifurcation is a Hopf bifurcation, which corresponds to the onset of moving rolls near the boundary. The onset of such rolls is the first step of one of the possible “routes to chaos” for shear layer flows.
1 Introduction
In this paper, we consider the incompressible Navier-Stokes equations in a half plane
$\partial_{t}u^{\nu}+(u^{\nu}\cdot\nabla)u^{\nu}-\nu\Delta u^{\nu}+\nabla p^{\nu}=f^{\nu},$ (1)
$\nabla\cdot u^{\nu}=0,$ (2)
together with the Dirichlet boundary condition
$u^{\nu}=0\qquad\hbox{for}\qquad y=0.$ (3)
We are interested in the nonlinear evolution of small perturbations of a given shear layer profile $U(y)=(U_{s}(y),0)$. Note that this shear layer profile is a stationary solution of Navier-Stokes equations provided we add the forcing term
$f^{\nu}=(-\nu\Delta U_{s},0).$
We assume that $U_{s}(y)$ is a smooth function, concave or convex, with $U_{s}(0)=0$, $\partial_{y}U_{s}(0)\neq 0$ and that $U_{s}$ converges exponentially fast at +$\infty$+ ∞ to some constant $U_{+}$.
Our results extend to the half-space, namely to three dimensions, without any change, all the flows being then invariant in $z$.
We can distinguish between two kinds of flows. The first possibility is that the shear flow is spectrally unstable for the linearized Euler equations. In this case it is nonlinearly unstable for Navier-Stokes equations. As shown in [6], small and smooth perturbations may grow and reach a magnitude $O(1)$ in $L^{\infty}$.
The second possibility is that the shear flow is spectrally stable for the linearized Euler equations. However, it is well-established in physics [4, 8], and proved in [5], that such shear flows are spectrally unstable for Navier-Stokes equations provided the viscosity is small enough. According to Rayleigh’s criterium, concave or convex flows belong to this second class. In order to fix the ideas, we focus on such flows, however our analysis extends to any flow of this second kind.
It is well known in physics than any shear layer is unstable and that small perturbations increase and ultimately lead to fully turbulent flows. According to [9] (page $403$), the development of instabilities may follow various scenarios, for instance the “spectral” scenario or the “by-pass” scenario.
In the “spectral” scenario, the unavoidable noise of any experiment triggers some linear instability, which first increases exponentially. Later, nonlinear mechanisms can no longer be neglected with respect to the linear growth. The linear instability get tamed by the nonlinear terms and saturates. This gives birth to a new “equilibrium” or new ”base state”, which in turn becomes unstable. Secondary instabilities appear and grow, leading to the destabilisation of this secondary base state, leading to a third base state. Instabilities then becomes more and more complex and close, leading to fully developed turbulence (so called “route to chaos”).
In this paper, we investigate the spectral scenario for particular shear layers, and in particular for the exponential profile
$U_{s}(y)=1-\exp(-\delta y),$ (4)
where $\delta$ is a positive constant. In physics, this shear layer is well-known to be linearly unstable provided the viscosity is small enough. The mathematical proof of this fact can be found in [5]. This justifies the first step of the “spectral” scenario, namely some initially small perturbations increase exponentially fast. In this article, we describe the second step of this scenario, namely the onset of a new base state: the exponentially growing perturbations get saturated by nonlinear effects and converge to a new equilibrium.
Let $\alpha$ be the horizontal Fourier number. Let us first recall that, provided $\nu$ is small enough, a shear flow is linearly unstable provided $\alpha$ lies in some interval $[\alpha_{-}(\nu),\alpha_{+}(\nu)]$, and stable outside this interval. For any $\alpha$ in this interval, there exists only one unstable eigenvalue $\lambda(\alpha,\nu)$ with corresponding eigenvector $\zeta(x,y)$ of the form
$\zeta(x,y)=\nabla^{\perp}\Bigl{[}e^{i\alpha x}\psi(y)\Bigr{]},$
for some stream function $\psi(y)$ depending on $\alpha$ and $\nu$. We are interested in the bifurcation for Fourier numbers close to $\alpha_{+}(\nu)$. We point out that for Fourier numbers close to $\alpha_{-}(\nu)$, resonances may appear due to the fact that integer multiples of $\alpha$ may belong to the instability region $[\alpha_{-}(\nu),\alpha_{+}(\nu)]$. We do not treat this situation here.
Let us now state our main results.
Theorem 1.1.
Let $\nu>0$ be small enough and $\alpha_{+}(\nu)$ such that $(\nu,\alpha_{+}(\nu))$ belongs to the upper marginally stable curve. Let $\zeta(x,y)$ be the eigenvector corresponding to the purely imaginary eigenvalue $\lambda(\alpha_{+}(\nu),\nu)$. Set
$\mu=\alpha-\alpha_{+}(\nu).$
Then, there exists a real number $\omega_{+}$ and complex numbers $c_{1}$ and $c_{3}$ such that for any $\mu$ sufficiently small, any sufficiently small bounded solution of (1)-(3) is of the form
$u^{\nu}(t,x,y)=U(y)+A(t)\zeta(x,y)+\bar{A}(t)\bar{\zeta}(x,y)+O\Bigl{(}|A|(|\mu|+|A|)\Bigr{)},$ (5)
where the complex-valued function $A$ satisfies the ordinary differential equation
${dA\over dt}=i\omega_{+}A+c_{1}\mu A+c_{3}A|A|^{2}+O\Bigl{(}|A|(|\mu|^{2}+|A|^{4})\Bigr{)}.$ (6)
Furthermore,
$i\omega_{+}=\lambda(\alpha_{+}(\nu),\nu),\qquad c_{1}=\partial_{\alpha}\lambda(\alpha_{+}(\nu),\nu).$
As usual in bifurcation theory, any small solution of (6) provides a solution of (1,2) of the form (5). The proof of this result relies on a center manifold reduction. Notice that the eigevalue $\lambda(\alpha_{+}(\nu),\nu)$ is purely imaginary, because $(\alpha_{+}(\nu),\nu)$ belongs to the marginally stable curve. Furthermore, $\Re c_{1}<0$, because $(\alpha_{+}(\nu),\nu)$ belongs to the upper marginally stable curve. There is no theoretical result on the coefficient $c_{3}$ and in particular on the sign of $\Re c_{3}$. However, the value of $c_{3}$ can be numerically evaluated [2], showing that $\Re c_{3}<0$. This leads to our second main result.
Theorem 1.2.
Under the hypotheses of Theorem 1.1, $\Re c_{1}<0$ and numerically, for the exponential profile (4), we have
$\Re c_{3}<0.$
Consequently, the exponential profile (4) undergoes a supercritical Hopf bifurcation on the upper marginal stability curve.
In this Hopf bifurcation, a stable time-periodic solution bifurcates for $\mu<0$ small enough, of the form
$u^{\nu}(t,x,y)=U(y)+|\mu|^{1/2}\Re\nabla^{\perp}\Bigl{[}\sqrt{c_{1r}\over c_{3r}}\,e^{i\omega_{+}t+i\alpha x}\psi(y)\Bigr{]}+O(|\mu|),$ (7)
where $c_{1r}=\Re c_{1}$ and $c_{3r}=\Re c_{3}$. In particular, the difference $u^{\nu}-U(y)$ is of size $O(\mu^{1/2})$ in $L^{\infty}$, and depends on $y$ and on the combination $i\omega_{+}t+i\alpha x$. It thus takes the form of a non-stationary “roll” in $(x,y)$ which moves with a constant speed $-\omega_{+}/\alpha$. The secondary base state is thus composed of moving rolls.
Up to the best of our knowledge, this result is the first bifurcation result for the genuine Navier-Stokes equations on a half space. In particular, it opens the way to the study of secondary instabilities.
2 Linear analysis
2.1 Orr-Sommerfeld equations
Let us first introduce the classical Orr-Sommerfeld equations. We refer to [1] for more details. Let $L_{NS}$ be the linearized Navier-Stokes operator near the shear layer profile $U$, namely
$L_{NS}\;v=(U\cdot\nabla)v+(v\cdot\nabla)U-\nu\Delta v+\nabla q,$ (8)
with $\nabla\cdot v=0$ and Dirichlet boundary condition. The solution of
$\partial_{t}v+L_{NS}\,v=f$ (9)
is then given by Dunford’s formula
[…]
in which $\Gamma$ is a contour on the “right” of the spectrum. This leads to study the resolvant of $L$
$(L_{NS}+\lambda)v=f,$ (11)
where $f$ is a given forcing term and $\lambda$ a complex number. Taking advantage of the divergence free condition, we introduce the stream function $\psi$ and take the Fourier transform in $x$ and the Laplace transform in $t$, which leads to look for solutions of the form
$v=\nabla^{\perp}\Bigl{(}e^{i\alpha(x-ct)}\psi(y)\Bigr{)}.$
Note that $\lambda=-i\alpha c$, according to the traditional notations. We also take the Fourier and Laplace transform of the forcing term $f$
$f=\Bigl{(}f_{1}(y),f_{2}(y)\Bigr{)}e^{i\alpha(x-ct)}.$
Taking the curl of (11) we then get the classical Orr-Sommerfeld equations
$Orr_{\lambda,\alpha,\nu}(\psi)=(U_{s}-c)(\partial_{y}^{2}-\alpha^{2})\psi-U_{s}^{\prime\prime}\psi-{\nu\over i\alpha}(\partial_{y}^{2}-\alpha^{2})^{2}\psi=i{\nabla\times f\over\alpha}$ (12)
where
×$\nabla\times(f_{1},f_{2})=i\alpha f_{2}-\partial_{y}f_{1}.$
The Dirichlet boundary condition gives
$\psi(0)=\partial_{y}\psi(0)=0.$ (13)
Note also that the “true” eigenvalue for linearised Navier-Stokes operator is $\lambda=-i\alpha c$, whereas $c$ is a “rescaled” eigenvalue, adapted for Orr-Sommerfeld equations. We will introduce
$\varepsilon=-{\nu\over i\alpha}.$
Note that the source term in the right hand side of (12) is a vorticity, whereas $\psi$ is a stream function. We will denote by $L_{NS}^{\alpha}$ the operator $L_{NS}$ after the Fourier transform in $x$.
2.2 Description of the spectrum of Orr-Sommerfeld operator
A complete description of the spectrum is obtained in [1]. It turns out that, for $\nu$ small enough, if $|\alpha|\lesssim\nu^{1/8}$, $L_{NS}$ has at most one eigenvalue $\lambda(\alpha,\nu)$ such that
$\Re\lambda(\alpha,\nu)\geq-\nu^{1/2}.$
Moreover, there exists $\alpha_{-}(\nu)$ and $\alpha_{+}(\nu)$ such that
$\alpha_{-}(\nu)\sim C_{-}\nu^{1/4},\qquad\alpha_{+}(\nu)\sim C_{+}\nu^{1/6}$
for some positive constants $C_{-}$ and $C_{+}$ and such that, if $\alpha_{-}(\nu)<\nu<\alpha_{+}(\nu)$ then $\Re\lambda(\alpha,\nu)>0$. Outside this interval, $\Re\lambda(\alpha,\nu)<0$. Moreover,
$\max_{\alpha}\Re\lambda(\alpha,\nu)\sim C_{0}\nu^{1/2}$ (14)
for some positive constant $C_{0}$. Similarly,
$\max_{\alpha_{+}(\nu)/2\leq\alpha\leq\alpha_{+}(\nu)}\Re\lambda(\alpha,\nu)\sim C_{1}\nu^{1/2}$ (15)
for some positive constant $C_{1}$.
The curve $\alpha_{+}(\nu)$ is called the “upper marginally stability branch” and $\alpha_{-}(\nu)$ the “lower marginally stability branch”.
This eigenvalue $\lambda(\alpha,\nu)$ is of multiplicity $1$. The associated eigenvector $\psi_{\lambda}$ can be expressed almost explicitly in terms of $U_{s}$ and of primitives of Airy functions. It is $C^{\infty}$ and exponentially decreasing, as well as all its derivatives.
The projection on the unstable eigenvector is given by the scalar product with the eigenvector of the adjoint of Orr-Sommerfeld operator. This eigenvector ${}^{t}\psi_{\lambda}$ is also almost explicitly given in terms of $U_{s}$ and of the $Ai$ Airy function.
2.3 Estimates on the resolvent
The Green function of Orr-Sommerfeld equation has been constructed and studied in details in [1] as $\nu$ goes to 00. In this paper however, $\alpha$, $c$ and $\nu$ are small but fixed, and thus can be considered as being of order $O(1)$. The inversion of (12) is then much easier.
Proposition 2.1.
Let $\psi$ and $f$ such that
$(\lambda+L_{NS}^{\alpha})\psi=f,$
and such that $\psi$ goes to 00 at infinity. Then for $\lambda$ large enough and for $\alpha$ close to $\alpha_{\star}$, we have
$\|\psi\|_{L^{2}}+\|\partial_{y}\psi\|_{L^{2}}+\|\partial_{y}^{2}\psi\|_{L^{2}}\lesssim|\lambda|^{-1}\|f\|_{L^{2}}.$ (16)
Proof.
We rewrite Orr-Sommerfeld equations as
$-\varepsilon\partial_{y}^{4}\psi-\tilde{c}\partial_{y}^{2}\psi+\alpha^{2}c\psi=E(\psi)$ (17)
where
$E(\psi)=-U_{s}\partial_{y}^{2}\psi+\alpha^{2}U_{s}\psi+U_{s}^{\prime\prime}\psi+\varepsilon\alpha^{4}\psi+f,$ (18)
where $f$ is some forcing term and where
$\tilde{c}=c-2\varepsilon\alpha^{2}.$
We recall that $\lambda=-i\alpha c$ and that $\alpha$ is fixed, thus large $\lambda$ correspond to large $c$ and $\tilde{c}$. We first consider the approximate equation
$-\varepsilon\partial_{y}^{4}\psi-\tilde{c}\partial_{y}^{2}\psi+\alpha^{2}c\psi=f.$ (19)
Looking for solutions of the form $e^{\mu y}$ leads to the characteristic equation
$-\varepsilon\mu^{4}-\tilde{c}\mu^{2}+\alpha^{2}c=0,$
and thus to
[…]
There exists thus two “slow solutions” $e^{\pm\mu_{s}y}$ where $\mu_{s}$ is of order $O(1)$ and two “fast solutions” $e^{\pm\mu_{f}y}$ where $\mu_{f}\sim\sqrt{\tilde{c}/\varepsilon}$, namely large when $\lambda$ is large.
The Green function of $G_{0}(x,y)$ (19), namely the solution corresponding to $f=\delta_{x}$ is the sum of an ”interior” Green function $G^{int}(x,y)$ which takes care of the source $f$, and of a ”boundary layer” Green function $G^{b}(x,y)$ which recovers the boundary condition.
We choose $G^{int}(x,y)$ of the form
$G^{int}(x,y)=ae^{-\mu_{f}|x-y|}+be^{-\mu_{s}|x-y|}$
for some constants $a$ and $b$ to be determined. Note that by construction, $G^{int}(x,y)$ and $\partial_{y}^{2}G^{int}(x,y)$ are continuous at $x$. We must write that $\partial_{y}G^{int}(x,y)$ is also continuous, and that $\partial_{y}^{3}G^{int}(x,y)$ has a jump of size $-1/\varepsilon$, which leads to
$\mu_{f}a+\mu_{s}b=0$
and
$\mu_{f}^{3}a+\mu_{s}^{3}b={1\over 2\varepsilon}.$
We thus have
$a={1\over 2\varepsilon}{1\over\mu_{f}}{1\over\mu_{f}^{2}-\mu_{s}^{2}}=O(|\lambda|^{-3/2})$
and
$b=-{1\over 2\varepsilon}{1\over\mu_{s}}{1\over\mu_{f}^{2}-\mu_{s}^{2}}=O(|\lambda|^{-1}).$
In particular, for large $\lambda$,
[…]
and similarly for integrals over $x$.
We now look for $G^{b}(x,y)$ under the form
$G^{b}(x,y)=a^{\prime}e^{-\mu_{f}y}+b^{\prime}e^{-\mu_{s}y}$
where the coefficients $a^{\prime}$ and $b^{\prime}$ depend on $x$. We then write
$G^{int}(x,0)+G^{b}(x,0)=\partial_{y}G^{int}(x,0)+\partial_{y}G^{b}(x,0)=0,$
which explicitly gives $a^{\prime}$ and $b^{\prime}$. It can then be checked that $G^{b}$ also satisfies an estimate of the form (20).
We now solve Orr-Sommerfeld, written under the form (18) by iteration, starting from
$\psi_{1}=G_{0}\star_{x}f$
where $G_{0}=G^{int}+G^{b}$. Note that $\|\psi_{1}\|_{L^{2}}\lesssim|\lambda|^{-1}\|f\|_{L^{2}}$. We then define by induction $E_{n}=E(\psi_{n})$ and $\psi_{n+1}=G_{0}\star_{x}E_{n}$. As $\|E_{n}\|_{L^{2}}\lesssim\|\psi_{n}\|_{L^{2}}$, this procedure converges, and the solution of (18) is given by
$\psi=\sum_{j}\psi_{j},$
and satisfies (16). ∎
3 Bifurcation analysis
We are interested in solutions which bifurcate close to the upper marginal stability branch $(\nu,\alpha_{+}(\nu))$. We fix $\nu$ small enough, so that the results in Section 2 hold, and take $\alpha$ as bifurcation parameter. It is also possible to consider $\nu$ as the bifurcation parameter, $\alpha$ being fixed.
3.1 Dynamical system formulation
For the system (1)-(3), we consider solutions of the form
$u^{\nu}(t,x,y)=U(y)+v(t,x,y),$
in which $U(y)$ is the shear profile, and we assume that the perturbation $v(t,x,y)$ is $2\pi/\alpha$-periodic in $x$. In order to avoid $\alpha$-dependent function spaces, we rescale variables by taking
$\tilde{x}=\alpha x,\qquad\tilde{y}=\alpha y,\qquad\tilde{t}=\alpha t,$ (21)
so that
$v(t,x,y)=\tilde{v}(\tilde{t},\tilde{x},\tilde{y}),$
is $2\pi$-periodic in $x$. Dropping the tilde, we obtain the system
$\partial_{t}v=-L_{NS}v-(v\cdot\nabla)v,$ (22)
$\nabla\cdot v=0,$ (23)
$v=0,\qquad\hbox{for}\qquad y=0,$ (24)
in which $L_{NS}$ is the scaled $L_{NS}$ operator
$L_{NS}\,v=-\Bigl{[}(U\cdot\nabla)v+(v\cdot\nabla)U-\alpha\nu\Delta v+\nabla q\Bigr{]}.$
For this system, we look for solutions which are $2\pi$-periodic in $x$.
As a first step, we formulate this system as a dynamical system
$\partial_{t}v=L_{\alpha}v+B(v),$ (25)
in a suitably chosen phase space ${\cal X}$, in which $L_{\alpha}$ is a closed operator with domain ${\cal Z}$, the phase space ${\cal X}$ includes the condition $\nabla\cdot v=0$ and the domain ${\cal Z}$ includes the remaining boundary conditions, and, more importantly, the gradient term $\nabla q$ is eliminated by a suitable projection. Such a formulation is well-known (see for instance [10], and [3] for the case of a cylindrical domain).
We take
${\cal X}=\Bigl{\{}v\in(L^{2}_{per}(\Omega))^{2}\quad|\quad\nabla\cdot v=0,\quad v_{2}|_{y=0}=0\Bigr{\}},$
where $\Omega={\mathbb{R}}\times(0,\infty)$, the subscript ${per}$ means that the functions are $2\pi$-periodic in $x$ and $v=(v_{1},v_{2})$. Next, we consider the Leray projection $\Pi_{0}:(L^{2}_{per}(\Omega))^{2}\to{\cal X}$, defined as the orthogonal projection on ${\cal X}$ with respect to the standard $(L^{2}_{per}(\Omega))^{2}$ scalar product, so that
$\Pi_{0}\nabla q=0.$
Then, projecting with $\Pi_{0}$, we obtain the dynamical system formulation of (25) with
$L_{\alpha}=-\Pi_{0}L_{NS},\qquad B(v)=-\Pi_{0}(v\cdot\nabla v).$
In this formulation, $L_{\alpha}$ is a closed linear operator in ${\cal X}$ with domain
${\cal Z}=\Bigl{\{}v\in(H^{2})^{2}\quad|\quad\nabla\cdot v=0,\quad v|_{y=0}=0\Bigr{\}}.$
Moreover, $B:{\cal Z}\to{\cal X}$ is well-defined and quadratic, thus $B$ is of class $C^{k}$ for any $k\geq 1$.
3.2 Center manifold reduction
The linear analysis in Section 2 allows us to apply a centre manifold theorem to the dynamical system (25) for $\alpha$ close to $\alpha_{+}(\nu)$, $\nu$ being fixed. This will prove the result in Theorem 1.1.
For notational convenience we write $\alpha_{+}=\alpha_{+}(\nu)$. The following properties of the linear operator $L_{\alpha_{+}}$ are an immediate consequence of the results in Section 2.
Lemma 3.1.
  • (i)
    The linear operator $L_{\alpha_{+}}$ has precisely two simple purely imaginary eigenvalues ±$\pm i\omega_{+}$ with associated eigenvectors $\zeta$ and $\bar{\zeta}$ of the form
    $\zeta(x,y)=e^{ix}\phi(y).$
    Furthermore, there exists $\sigma>0$ such that the spectrum $\sigma(L_{\alpha_{+}})$ satisfies
    $\sigma(L_{\alpha_{+}})\setminus\{\pm i\omega_{+}\}\subset\{\lambda\in{\mathbb{C}}\quad|\quad\Re\lambda<-\sigma\}.$
  • (ii)
    There exist constants $C>0$ and $\omega_{0}>0$ such that the following estimate holds
    $\|(L_{\alpha_{+}}-i\omega)^{-1}\|_{{\cal X}\to{\cal X}}\leq{C\over|\omega|},$
    for any $|\omega|\geq\omega_{0}$.
We set $\alpha=\alpha_{+}+\mu$, so that $\mu$ is now a small bifurcation parameter. We rewrite the dynamical system (25) as
$\partial_{t}v=L_{+}v+L_{\mu}v+B(v),$ (26)
with
$L_{+}=L_{\alpha_{+}},\qquad L_{\mu}=L_{\alpha_{+}+\mu}-L_{\alpha_{+}}=\mu\nu\Delta.$
The properties in Lemma 3.1 and the property that $B$ is quadratic imply that the hypotheses of the centre manifold theorem [7, Theorem $2.20$, Chapter $2$] are satisfied, for $\mu$ sufficiently small. As a consequence, small bounded solutions of (26) are of the form
$v(t)=A(t)\zeta+\bar{A}(t)\bar{\zeta}+O\Bigl{(}|A|(|\mu|+|A|)\Bigr{)},$
in which $\zeta$ and $\bar{\zeta}$ are the eigenvectors of $L_{+}=L_{\alpha_{+}}$ in Lemma 3.1 and $A$ is a complex-valued function satisfying an ordinary differential equation of the form
${dA\over dt}=i\omega_{+}A_{+}+f(\mu,A,\bar{A}).$ (27)
Note that $\bar{A}(t)$ satisfies the complex conjugated equation.
We can obtain a more precise description of $f(\mu,A,\bar{A})$ by taking into account the invariance of the equations under translations in $x$. As a consequence, the system (26) is equivariant under the action of the continuous group
$\tau_{a}v(x,y)=v(x+a,y),\qquad\forall a\in{\mathbb{R}}.$
This symmetry is inherited by the reduced system (27). Since
$\tau_{a}\zeta=e^{ia}\zeta,\qquad\forall a\in{\mathbb{R}},$
the action of $\tau_{a}$ on $A(t)$ reads
$\tau_{a}A=e^{ia}A,\qquad\forall a\in{\mathbb{R}}.$
Consequently, the equivariance of (27) under the action of $\tau_{a}$ implies that
$f(e^{ia}A,e^{-ia}\bar{A},\mu)=e^{ia}f(A,\bar{A},\mu),$
so that
$f(A,\bar{A},\mu)=Ag(|A|^{2},\mu),$
according to [7, Lemma $2.4$, Chapter $1$]. To leading order, we find the equation
${dA\over dt}=i\omega_{+}A+c_{1}\mu A+c_{3}A|A|^{2},$ (28)
the higher order terms being $O(|A|(|\mu|^{2}+|A|^{4}))$. Taking into account the scaling (21) this proves the result in Theorem 1.1 (in which $\omega_{+}$, $c_{1}$ and $c_{3}$ are the ones above multiplied by $\alpha_{+}$).
3.3 Hopf bifurcation
To analyze the solutions of (28), we need to compute the signs of the real parts of the coefficients $c_{1}$ and $c_{3}$.
The coefficient $c_{1}$ can be related to the eigenvalue $\lambda(\alpha,\nu)$. Indeed, $\lambda(\alpha,\nu)$ is the eigenvalue of the linearization of the reduced equation (27) at $A=0$ multiplied by $\alpha$, due to the rescaling of $t$. This implies that
$\lambda(\alpha_{+}+\mu,\nu)=(\alpha_{+}+\mu)\Bigl{(}i\omega_{+}+c_{1}\mu+O(\mu^{2})\Bigr{)},$
which then allows to determine the sign of $\Re c_{1}$. As the flow is stable for $\alpha>\alpha_{+}(\nu)$, we have that $\Re\lambda(\alpha_{+}+\mu,\nu)<0$ when $\mu>0$. This implies that
$\Re c_{1}<0,$
because $\partial_{\alpha}\Re\lambda(\alpha_{+},\nu)\not=0$.
To compute $c_{3}$, one can use the method given in [7, Chapter $1$, page $16$] which provides an analytical formula of $c_{3}$. In the present case this type of computation seems out of reach, because it requires an explicit inversion of the Orr-Sommerfeld equation. However, it is possible to evaluate $c_{3}$ numerically [2]. In the case of the exponential profile, this numerical computation shows that
$\Re c_{3}<0.$
The signs of the coefficients $c_{1}$ and $c_{3}$ found above show that for the differential equation (28) a supercritical Hopf bifurcation occurs at $\mu=0$, and the same property holds for the reduced equation (27) (e.g., see [7, Theorem 2.6, Chapter 1]). More precisely, we have the following properties:
  • (i)
    for any sufficiently small $\mu$, the differential equation (27) possesses the equilibrium $A=0$ which is stable if $\mu>0$ and unstable if $\mu<0$;
  • (ii)
    for any sufficiently small $\mu<0$, the differential equation (27) possesses a unique periodic orbit $A_{\mu}=O(|\mu|^{1/2})$ which is stable.
Furthermore, setting $c_{1r}=\Re c_{1}$ and $c_{3r}=\Re c_{3}$, we have that
$|A_{\mu}(t)|=\sqrt{\frac{c_{1r}}{c_{3r}}}|\mu|^{1/2}+O(|\mu|).$
Going back to the dynamical system (25) we obtain the stable periodic solution
$v(t)=2\Re(A_{\mu}(t)\zeta)+O(|\mu|)=2|\mu|^{1/2}\sqrt{\frac{c_{1r}}{c_{3r}}}\,\Re\left(e^{i(x+\omega_{+}t)}\phi(y)\right)+O(|\mu|).$
for $\mu<0$ sufficiently small, and going back to the original system (1)-(3) we obtain the solution (7).
Acknoledgements
D. Bian is supported by NSFC under the contract 12271032.
Conflict of interest
The authors state that there is no conflict of interest.
Data availability
Data are not involved in this research paper.